# Subradiant Decay in 2D and 3D Atomic Arrays

### v1

@@ I Introduction
In this context, infinite lattices offer interesting insights on the scattering properties of these arrays. The single-excitation eigenmodes of the scattering problem are translationally invariant and obey Bloch’s theorem. This allows one to define the quasi-momentum $\mathbf{k}$ and calculate explicitly the collective frequency shift and decay rate of each of these modes [15]. Differently, in finite systems these collective features can only be accessed by numerical diagonalization, leading to severe limitations on the system size which can be simulated. Furthermore, most studies on subradiance in ordered systems have focused on one-dimensional (1D) atomic chains [6, 7, 8, 17, 18, 19, 20, 21] and two-dimensional (2D) arrays [9, 14, 15, 22, 13], with investigations on the three-dimensional configurations lacking [23]. We note that while the band structure of 2D and 3D arrays have been studied through the density of states [24, 25, 26], the connection between this phenomenon and subradiance has not been yet clarified.
@@ II.1 Coupled Dipole Model
The single-excitation states are $|j\rangle=|g_{1},\dots,e_{j},\dots,g_{N}\rangle$ and $\mathbf{k}\in{\cal S}$, where $\cal S$} is the volume of first Brillouin zone in the reciprocal lattice, i.e., $|k_{x}|,|k_{y}|,|k_{z}|\leq\pi/d$. It includes the symmetric Dicke state [1] for $\mathbf{k}=0$ , as well as the timed-Dicke state introduced in Ref. [31, 32] for a driving with wavevector $\mathbf{k}=\mathbf{k}_{0}$. The family $\{|\mathbf{k}\rangle\}$ satisfies the completeness relation
@@ II.3 Collective decay rate
Eq. (10) corresponds to the Fourier transform of $\Gamma_{jm}$ only if the lattice is infinite, when modes (7) are the Bloch states and set $\{\mathbf{k}\}$ is discrete. Nevertheless, treating $\mathbf{k}$ as a continuous variable in Eq. (10) is a good approximation for the mode when $N$ is sufficiently large [15]. We hereafter refer to $\Gamma(\mathbf{k})$ as the “continuous spectrum” of the decay rate. We highlight that we are not referring to the discrete spectrum of the eigenvalues of the system, described by the non-Hermitian Hamiltonian (1), but rather to the imaginary part of the expectation value of the effective Hamiltonian on the collective state of Eq. (7), describing the interference among the (single-photon) fields emitted by $N$ atoms. In particular, in the infinite $N$ limit, $\Gamma(\mathbf{k})$ is precisely the decay rate of mode $\mathbf{k}$, yet in the finite-$N$ case, the exact rate must be obtained by diagonalizing the full matrix (1).
@@ III.1 General expression of the decay rate
$\times$ $\mathrm{sinc}^{2}\left[\left(k_{x}-\frac{2\pi m_{x}}{d}-k_{0}\sin\theta_{0}\cos\phi_{0}\right)\frac{dN_{x}}{2}\right]$
$\times$ $\mathrm{sinc}^{2}\left[\left(k_{y}-\frac{2\pi m_{y}}{d}-k_{0}\sin\theta_{0}\sin\phi_{0}\right)\frac{dN_{y}}{2}\right].$
@@ III.2 Infinite square array
with $\mathbf{g}=(g_{x},g_{y})$ the reciprocal array vector introduced previously. Fig. 1 shows $\Gamma(k_{x},k_{y})/\Gamma_{0}$ as a function of $k_{x}d$ and $k_{y}d$ for $d=\lambda_{0}/5$ and $d=\lambda_{0}$ ($\lambda_{0}=2\pi/k_{0}$ is the atomic transition wavelength), and for dipoles oriented along the $x$-axis. The black areas correspond to dark states, with $\Gamma(k_{x},k_{y})=0$: They have a suppressed emission, akin to subradiant modes, yet reaching $\Gamma(k_{x},k_{y})=0$ in this infinite-size limit. Note that for $d<\lambda_{0}/2$ only the terms $g_{x}=0$ and $g_{y}=0$ contribute to the sums in Eq. (LABEL:Gamma:2D:infinite), while for $\lambda_{0}/2<d<\lambda_{0}$, the terms $g_{x}=\pm 2\pi/d$ and $m_{y}=\pm 2\pi/d$ also contribute.
@@ V Conclusion
While the present work addressed the single-excitation regime, the possibility to use two-level atoms to manipulate multiple-photons states calls for the extension of such methods to the many-body regime [33, 34, 35]. In particular, combining the long lifetimes of subradiant modes with the photon-photon interactions achieved through dipolar or Rydberg interactions could allow for the creation of tunable atomic platforms for photon manipulation.

### v2

@@ I Introduction
In this context, infinite lattices offer interesting insights on the scattering properties of these arrays. The single-excitation eigenmodes of the scattering problem are translationally invariant and obey Bloch’s theorem. This allows one to define the quasi-momentum $\mathbf{k}$ and calculate explicitly the collective frequency shift and decay rate of each of these modes [15]. Differently, in finite systems these collective features can only be accessed by numerical diagonalization, leading to severe limitations on the system size which can be simulated. Furthermore, most studies on subradiance in ordered systems have focused on one-dimensional (1D) atomic chains [6, 7, 8, 17, 18, 19, 20, 21] and two-dimensional (2D) arrays [9, 14, 15, 22, 13], with three-dimensional configurations less investigated  [23, 24, 25]. We note that while the band structure of 2D and 3D arrays have been studied through the density of states [26, 27, 28], the connection between this phenomenon and subradiance has not been yet clarified.
@@ II.1 Coupled Dipole Model
The single-excitation states are $|j\rangle=|g_{1},\dots,e_{j},\dots,g_{N}\rangle$ and $\mathbf{k}\in{\cal S}$, where $\cal S$} is the volume of first Brillouin zone in the reciprocal lattice, i.e., $|k_{x}|,|k_{y}|,|k_{z}|\leq\pi/d$. It includes the symmetric Dicke state [1] for $\mathbf{k}=0$ , as well as the timed-Dicke state introduced in Ref. [33, 34] for a driving with wavevector $\mathbf{k}=\mathbf{k}_{0}$. The set $\{|\mathbf{k}\rangle\}$ satisfies the completeness relation
@@ II.3 Collective decay rate
Eq. (10) corresponds to the Fourier transform of $\Gamma_{jm}$ only if the lattice is infinite, when modes (7) are the Bloch states and set $\{|\mathbf{k}\rangle\}$ is discrete. Nevertheless, treating $\mathbf{k}$ as a continuous variable in Eq. (10) is a good approximation for the mode when $N$ is sufficiently large [15]. We hereafter refer to $\Gamma(\mathbf{k})$ as the “continuous spectrum” of the decay rate. We highlight that we are not referring to the discrete spectrum of the eigenvalues of the system, described by the non-Hermitian Hamiltonian (1), but rather to the imaginary part of the expectation value of the effective Hamiltonian on the collective state of Eq. (7), describing the interference among the (single-photon) fields emitted by $N$ atoms. In particular, in the infinite $N$ limit, $\Gamma(\mathbf{k})$ is precisely the decay rate of mode $\mathbf{k}$, yet in the finite-$N$ case, the exact rate must be obtained by diagonalizing the full matrix (1).
@@ III.1 General expression of the decay rate
$\times$ $\mathrm{sinc}^{2}\left[\left(k_{x}-g_{x}-k_{0}\sin\theta_{0}\cos\phi_{0}\right)\frac{dN_{x}}{2}\right]$
$\times$ $\mathrm{sinc}^{2}\left[\left(k_{y}-g_{y}-k_{0}\sin\theta_{0}\sin\phi_{0}\right)\frac{dN_{y}}{2}\right].$
@@ III.2 Infinite square array
with $\mathbf{g}=(g_{x},g_{y})$ the reciprocal array vector introduced previously. Fig. 1 shows $\Gamma(k_{x},k_{y})/\Gamma_{0}$ as a function of $k_{x}d$ and $k_{y}d$ for $d=\lambda_{0}/5$ and $d=\lambda_{0}$ ($\lambda_{0}=2\pi/k_{0}$ is the atomic transition wavelength), and for dipoles oriented along the $x$-axis. The black areas correspond to dark states, with $\Gamma(k_{x},k_{y})=0$: They have a suppressed emission, akin to subradiant modes, yet reaching $\Gamma(k_{x},k_{y})=0$ in this infinite-size limit. Note that for $d<\lambda_{0}/2$ only the terms $g_{x}=0$ and $g_{y}=0$ contribute to the sums in Eq. (LABEL:Gamma:2D:infinite), while for $\lambda_{0}/2<d<\lambda_{0}$, the terms $g_{x}=\pm 2\pi/d$ and $g_{y}=\pm 2\pi/d$ also contribute.
@@ V Discussion
Another important aspect is the preparation of the generalized Dicke state $|\mathbf{k}\rangle$ of Eq.(7). For example, in Ref. [22] the state $\mathbf{k}=0$ has been excited by a laser perpendicular to a 2D atomic array with $k_{0}d\approx 4.27$, such that weak subradiance has been observed (see Fig.2b). More generally, a possible strategy to address selectively subradiant modes ($k_{x}>k_{0}$) in a 2D lattice with $k_{0}d<\pi$ could be to set the array close to a dielectric or a metallic substrate and exploit the evanescent field generated in the total internal reflection or in surface plasmonic wave. Similar approaches have been investigated in Ref. [35, 36]. An alternative is to drive a localized section of the array, which will excite many modes at the same time: At later time, only the subradiant modes will survive with a substantial population [37]. However, very few photons are emitted through this protocol (subradiant modes get weakly populated only), making this configuration challenging experimentally.
Finally, while the present work addresses the single-excitation regime, the technique proposed can in principle be extented to states with multiple excitations. More specifically, the possibility to use two-level atoms to create and manipulate multiple-photon states calls for the extension of single-photon results to the many-body regime [38, 39, 40]. In particular, combining the long lifetimes of subradiant modes with the photon-photon interactions achieved through dipolar or Rydberg interactions could allow for the creation of tunable atomic platforms for photon manipulation.